OpenStudy (anonymous):

Find x − 3y if x = −3 and y = −2 . I need help.

OpenStudy (pfenn1):

They are asking you to solve the expression when you replace the x and y in the expression with -3 and -2. Like this.\[x − 3y =(-3)-3(-2)=?\]

OpenStudy (anonymous):

yess i know that part . but i always mess my negative and positive signs up when i do it . i forget the rules .

OpenStudy (pfenn1):

-(-3)=+3 -3(-2)=+6 Does that help?

OpenStudy (anonymous):

yes it did (: thanks